OpenStudy (anonymous):

use an associative law to find an expression equivalent to b+(a+75)

OpenStudy (anonymous):

(b+a)+75

Parth (parthkohli):

Associative law says: \(\Large \color{MidnightBlue}{\Rightarrow (a + b) + c = a + (b + c) }\) Similarly, \(\Large \color{MidnightBlue}{\Rightarrow b + (a + 75) = (b + a) + 75 }\)

OpenStudy (anonymous):

that is wht i wrote n it kept sayin it wrong

OpenStudy (anonymous):

excuse me, can i write it the same way for t+(s+50)